# Large Language Models in Argument Mining: A Survey

### v2

@@ 1 Introduction
However, despite this breadth, most existing reviews remain anchored in the pre-LLM era. They emphasize supervised learning and feature engineering, formal annotation schemas, or segmented task pipelines, and so far have only cursorily addressed the revolutionary role of pre-trained language models. In response, a new wave of empirical literature has emerged—exemplified by work on LLM-driven argument summarization, argument extraction, and relationship classification (Chen et al., 2024; Altemeyer et al., 2025; Guida et al., 2025). Yet, there is still no review that systematically integrates these advances, examines how LLMs reshape the boundaries and synergies of AM subtasks, and reconnects these technical developments to foundational argumentation theory and annotation practice.
In response, this survey offers a timely and theory-informed synthesis of how modern Large Language Models are redefining argument mining. We present AM not as a linear sequence of isolated subtasks but as a network of interdependent LLM-powered components, united by prompt engineering, chain-of-thought reasoning, few-shot learning, and retrieval-augmented generation. We examine these innovations across a wide range of domains—from legal and scientific text to social media, education, peer reviews, and emerging multi-modal content—surveying benchmark results from the past three years as well as emphasizing the empirical gaps in long-context reasoning, multilingual adaptation, model interpretability, and efficient deployment.
Building on this foundation, the rest of the survey unfolds as follows.Section 2 catalogs datasets and benchmarks—emphasizing recent large, cross-domain, and multimodal resources suited for LLM evaluation. Section 3 develops a taxonomy of key AM subtasks—claim detection, premise identification, relation parsing, quality assessment, and summarization—exploring how LLM techniques like prompting, chain-of-thought, few-shot learning, and retrieval-augmented generation reshape each. Section 4 introduces model architectures and key LLM-era techniques, detailing encoder–decoder designs, prompting strategies, fine-tuning approaches, and multimodal integration relevant to argument mining. Section 5 outlines evaluation methodologies, including metrics, human-in-the-loop assessments, and benchmarking practices tailored to AM. Section 6 delves into ongoing challenges—such as long-context reasoning, domain transfer, interpretability, annotation bottlenecks, and computational efficiency—while highlighting emerging research trends. Finally, Section 7 presents our conclusions and proposes a forward-looking roadmap for advancing LLM-powered computational argumentation.
2 Dataset and Benchmark
The landscape of resources for argument mining has undergone a significant evolution, marked by a clear trajectory from the analysis of static, decontextualized text towards a more nuanced understanding of argumentation in its natural settings. This progression is evidenced by the creation of novel datasets, particularly between 2021 and 2024, which not only expand the volume of available data but also fundamentally advance the complexity of tasks and the scope of inquiry within the field. These resources increasingly prioritize the dynamic, interactive, and multimodal facets of argumentation, reflecting the field’s growing maturity.
2.1 Capturing Argumentation in Dialogue
A primary vector of recent dataset development has been the focus on conversational argumentation. Recognizing that real-world arguments are predominantly forged through interaction, recent efforts have sought to capture the intricate dynamics of dialogue. A significant development in this area is the ADEA (Argumentative Dialogue Dataset on Ethical Issues Concerning Future A.I. Applications) Hauptmann et al. (2024). This dataset provides dialogues centered on the complex and value-laden topic of AI ethics, enabling research into how arguments are proposed, contested, and defended within a specialized, high-stakes domain.
Moving towards a more granular analysis of dialogue mechanics, the CAMAL (Conversational Argument Move Analysis) dataset introduces a multi-label framework for annotating "conversational moves" Lai et al. (2024). This approach moves beyond the traditional support/attack dichotomy to capture the multifaceted functions of utterances, such as questioning, clarifying, or challenging a premise. Such a fine-grained annotation scheme more accurately reflects the fluidity of natural dialogue and supports the development of models for more sophisticated dialogical systems. The challenge of parsing multi-party dialogue is addressed by the work of Mao et al., who developed a new argumentative dataset alongside a hierarchical graph fusion network Mao et al. (2024). Their contribution underscores the necessity of understanding global discourse structure for accurately resolving local argumentative relations in complex, multi-participant settings. Reinforcing this theme, the work of Mezza et al. demonstrates that exploiting dialogue acts and contextual information significantly improves the identification of argumentative relations in online debates Mezza et al. (2024).
2.2 Multi-modality Benchmark
As the number of specialized datasets has grown, the issue of fragmentation has become more salient, prompting efforts to foster model generalization and broaden the modalities of analysis. The ARIES (A General Benchmark for Argument Relation Identification) project represents a critical step towards unification Gemechu et al. (2024). By systematically integrating multiple existing datasets from diverse domains and annotation traditions, ARIES establishes a standardized benchmark for a core argument mining task. This initiative facilitates more equitable model comparisons and encourages the development of more robust algorithms that are not over-fitted to the idiosyncrasies of a single corpus. The importance of data composition is further explored by Schiller et al., who find that for topic-dependent argument mining, dataset diversity often has a greater impact on model performance than sheer size, providing crucial guidance for future data collection efforts Schiller et al. (2024).
The field’s scope has also expanded beyond purely textual data. The M-Arg dataset was a pioneering contribution in this direction, offering a multimodal resource for argument mining in political debates that includes both audio recordings and aligned transcripts Mestre et al. (2021). This enables the joint analysis of linguistic content and acoustic-prosodic features (e.g., pitch, tone, pauses), allowing researchers to investigate how non-verbal cues contribute to the expression of argumentative functions. M-Arg thus provides the empirical foundation for building computational models capable of a more holistic understanding of argumentation.
2.3 Argument Structure
Concurrent with the expansion of scope, there is a clear trend towards deepening the analysis of core argumentative concepts and applying them in highly specialized domains. The legal field, with its inherently structured and argumentative nature, remains a key area of application. The dataset of annotated European Court of Human Rights (ECHR) cases presented by Mumford et al. offers detailed insights into the intricacies of legal reasoning Mumford et al. (2024). Similarly, in the medical domain, the CasiMedicos-Arg dataset annotates explanatory argumentative structures within medical question-answering contexts, linking argument mining directly to the critical need for explainability in clinical AI Sviridova et al. (2024a, b).
The assessment of argument quality has emerged as another major research frontier. Large-scale datasets such as ArgAnalysis35K Joshi et al. (2023) and the work by Ivanova et al. on quality dimensions Ivanova et al. (2024) signal a shift from identifying the presence of an argument to evaluating its intrinsic merit. These resources facilitate the development of models that can assess arguments along multiple axes, including logic, sufficiency, and clarity, which is essential for applications such as writing support and automated feedback systems. Furthermore, there is a renewed focus on the foundational logical and philosophical underpinnings of argumentation. The EthiX dataset for argument scheme classification in ethical debates exemplifies this, connecting computational analysis back to established argumentation theory Bezou-Vrakatseli et al. (2024). Similarly, the creation of a dataset for causality identification by Al-Khatib et al. pushes the field towards identifying the deeper reasoning patterns within argumentative texts Khatib et al. (2023).
2.4 Data Creation and Evaluation in LLM Era
The advent of Large Language Models (LLMs) has introduced a new dynamic into the creation and use of argumentative datasets. On one hand, LLMs are being leveraged as tools to overcome data scarcity. The work of de Souza and Dragoni, for instance, explores the use of LLMs to augment persuasive argument datasets, offering a scalable method for generating new training instances de Souza and Dragoni (2024). On the other hand, the argumentative capabilities of LLMs are themselves becoming a central object of study. Researchers are now creating datasets specifically designed to probe and evaluate the quality of arguments generated by models like ChatGPT. The work by Rocha et al. provides a methodology and dataset to assess "good, bad, and ugly" machine-generated arguments Rocha et al. (2024), while Djouvas et al. present a comparative study and dataset for argument strength classification that critically evaluates the performance of LLMs against other models Djouvas et al. (2024). This reflexive turn, where LLMs are used both to create and to be evaluated by data, marks a significant new phase in the field, enabling a more nuanced understanding of the strengths and limitations of current generative models in the context of complex human reasoning.
@@ 3.2.4 Multil-modal Evidence Detection
Multi-modal rumor detection has seen significant enhancements through dual evidence mechanisms. Huang et al. Huang et al. (2025) proposed dual evidence enhancement alongside text-image similarity awareness, significantly improving misinformation detection efficacy. Similarly, Wu and Cao Wu and Cao (2024) developed an evidence-aware multimodal framework tailored specifically for Chinese social media, effectively capturing the nuances inherent in multimodal data.
@@ 3.3 Stance Detection
Beyond generic domains, stance detection has been used in medical and health misinformation contexts. Davydova et al. (2024) analyze COVID-19 tweet stances; Sun et al. (2024) study HPV vaccine stance detection using in-context learning and fine-tuning. Akash et al. (2024) assess the feasibility of open-target stance detection using LLMs.

### v3

@@ 1 Introduction
However, despite this breadth, most existing reviews remain anchored in the pre-LLM era. They emphasize supervised learning and feature engineering, formal annotation schemas, or segmented task pipelines, and so far have only cursorily addressed the revolutionary role of pre-trained language models. In response, a new wave of empirical literature has emerged—exemplified by work on LLM-driven argument summarization, argument extraction, and relationship classification (Chen et al., 2024; Altemeyer et al., 2025; Guida et al., 2025). Yet, there is still no review that systematically integrates these advances, examines how LLMs reshape the boundaries and synergies of AM subtasks, examining how advances in one area reshape assumptions in another, and reconnects these technical developments to foundational argumentation theory and annotation practice. We treat argument mining as an interlocking system whose components—tasks, data, metrics, and models—evolve together under the pressure of foundation-model capabilities.
This survey aims to fill that gap. We treat argument mining as an interlocking system whose components, tasks, data, metrics, and models, evolve together under the pressure of foundation-model capabilities. Our review makes three contributions: (i) Task re-appraisal. Chapter 3 revisits the canonical AM sub-tasks—claim and evidence detection, relation parsing, stance, argument quality, and summarisation showing how prompting, chain-of-thought, and retrieval-augmented generation blur established boundaries. (ii) Data and evaluation synthesis. Chapters 4 and 5 map the co-evolution of datasets and metrics, highlighting how LLM-assisted annotation and soft-label evaluation change what “state-of-the-art” means. (iii) Technology survey and agenda-setting. Chapter 6 distils common architectural patterns for LLM-era AM (instruction tuning, role-prompt cascades, graph-augmented decoders), while Chapter 7 identifies pressing open problems: long-context reasoning, multimodal and multilingual robustness, bias auditing for machine-generated data, and cost-efficient deployment.
2 Methodology
To obtain a comprehensive, up-to-date picture of research on argument mining (AM) in the large-language-model era, we followed a three-step protocol inspired by prior survey work (Ivanova et al., 2024). Figure 1 gives an overview.
Sources. We first manually queried two bibliographic repositories that jointly cover virtually all peer-review venues in computational linguistics and artificial-intelligence research: (i) the Digital Bibliography and Library Project [^1] https://dblp.org (DBLP) and (ii) the ACL Anthology [^2] https://aclanthology.org. Because the terminology used by the field is heterogeneous, we constructed a Boolean query that targets both generic and fine-grained AM vocabulary (e.g. argument, claim, premise, evidence, key point) and action like mining, extraction, detection. Following recent survey conventions, we limited the search to the last ten calendar years (2015–2025, inclusive), specific for recent 4 years of LLMs era. We find around 210 unique publications that match the query and fall into the 10-year window. We de-duplicated pre-prints against their camera-ready versions by DOI.
Dataset
Many AM papers introduce new corpora or benchmarks that subsequently shape the evaluation landscape. We therefore flagged all items whose title or booktitle fields contained any of the strings dataset, corpus, benchmark, collection, or shared task. This yielded around 70 dataset papers, i. e., works that either (a) release an original annotated resource or (b) substantially extend an existing one. In this work we focus on 40 of represent work to discuss.
Iterative Snowballing
To mitigate false negatives, especially papers that use idiosyncratic task names (e.g. “key-point analysis”, “argument quality ranking”) we employed one rounds of backward and forward snowballing on the seed set’s references. The procedure added 38 further works (6.9 % of the corpus), all of which passed the same year and scope filters. Because every snowballed item also appears in DBLP or the ACL Anthology, the final corpus size is 250 research papers.
Figure 1: Paper-collection workflow and resulting dataset.
@@ 3.2.4 Multil-modal Evidence Detection
Multi-modal rumor detection has seen significant enhancements through dual evidence mechanisms. Huang et al. (2025) proposed dual evidence enhancement alongside text-image similarity awareness, significantly improving misinformation detection efficacy. Similarly, Wu and Cao Wu and Cao (2024) developed an evidence-aware multimodal framework tailored specifically for Chinese social media, effectively capturing the nuances inherent in multimodal data.
@@ 3.2.5 Counter-Evidence Detection
3.3 Relation Identification
If claim detection marks the nodes of an argumentative graph, relation identification supplies its edges. Early work treated the task as sentence-pair classification—“support”, “attack”, or “unrelated”—in tightly controlled genres such as German news commentary (Eckle-Kohler et al., 2015). Genre expansion quickly revealed edge sparsity and annotation drift: the AMPERSAND corpus of Reddit threads, for example, required additional labels like rebuttal and mitigation to capture conversational nuance (Chakrabarty et al., 2020). Methodologically, the field has moved from feature-rich SVMs to graph-neural encoders that propagate information across discourse trees (DBLP:conf/acl/ZhangNL23). Yet even state-of-the-art models stumble when relations span paragraphs, rely on implicit premises, or cross language boundaries. .
3.4 Argument-Quality Assessment
Beyond structural edges lies the question of how good an argument actually is. Early datasets split quality into separate facets—logical sufficiency (Stab and Gurevych, 2017), rhetorical convincingness (Habernal and Gurevych, 2016), or holistic five-point scales (Lauscher et al., 2020). Recent work combines these dimensions and asks models to assign soft labels that reflect annotator disagreement (Plank, 2022). Large language models have further blurred boundaries: prompt-based “LLM judges” now produce zero-shot sufficiency scores that rival supervised baselines, but also inherit topic and sentiment biases (Chen and Eger, 2025). Despite impressive headline gains, fine-grained error analyses show that LLMs still favour fluent but logically thin arguments over less polished yet better-supported ones.
@@ 3.5 Stance Detection
Beyond generic domains, stance detection has been used in medical and health misinformation contexts. Davydova et al. (2024) analyze COVID-19 tweet stances; Sun et al. (2024a) study HPV vaccine stance detection using in-context learning and fine-tuning. Akash et al. (2024) assess the feasibility of open-target stance detection using LLMs.